OpenStudy (anonymous):

prove: cot^2(w)/csc(w) -1 = csc(w) +1

OpenStudy (anonymous):

cot^2 = csc^2-1 csc^2-1/csc-1 = csc+1 (csc+1)(csc-1)/csc-1 = csc+1

OpenStudy (anonymous):

kk got it from here, thank you